# Scaling Laws for Predicting Downstream Performance in LLMs

### v1

@@ Appendix B Analytical Form to Fit FLOPs-to-Performance Curve
Appendix C Compare with Llama-3 Approach
We compare with the Llama-3 approach for downstream task prediction (Dubey et al., 2024). They suggest initially estimating the negative log-likelihood (NLL) of the target answer based on the computational cost in FLOPs, followed by using this NLL to model the task performance through a sigmoid function. The comparison results are shown in Fig. 11. We find that the two-stage approach proposed in Dubey et al. (2024) fails to effectively estimate the performance based on our data points, compared to FLP.
Appendix D MMLU Experiment
Figure 12: The performance prediction on MMLU using FLP.
Appendix E FLP-M: Fit Curve for Ablation Study
@@
The FLOPs-to-Loss fit curves are in Fig. 13 and the Loss-to-Performance fit curves are in Fig. 14. We observe that M4 in Tab. 3 offers more stable and accurate predictions for domain-specific loss, with the combined approach of domain loss and neural network estimation delivering the best overall downstream performance prediction.
@@
Appendix F Using Domain Loss in FLP
We explore the application of FLP-M during pre-training on a consistent distribution (the experimental setting described in §4), and compare it with FLP. The fitting curves are shown in Fig. 15 and the results of relative prediction error are shown in Fig. 16. We show that FLP-M fails to effectively predict the performance of target LLMs when sampling LMs are pre-trained on a fixed distribution. This ineffectiveness is attributed to the closely related domain-specific validation losses among the sampling LMs within the same training distribution, which suggests that decomposing the pre-training validation loss yields no additional information in this pre-training setting. Thus, estimating five domain-specific loss, rather than a single average validation loss, can further increase the risk of error propagation. Moreover, using highly correlated features as neural network inputs may lead to overfitting.

### v2

@@ Appendix C Analytical Form to Fit FLOPs-to-Performance Curve
Appendix D Compare with Llama-3 Approach
We compare with the Llama-3 approach for downstream task prediction (Dubey et al., 2024). Our implementation strictly adheres to the two-stage performance prediction framework outlined in the Llama-3 paper: (1) Stage 1: We estimate the negative log-likelihood (NLL) based on FLOPs using the prescribed analytical forms; (2) Stage 2: We map the estimated NLL to task performance through a sigmoid function. The comparison results are shown in Fig. 12. We find that the two-stage approach proposed in Dubey et al. (2024) fails to effectively estimate the performance based on our data points, compared to FLP. The divergence between our results and those reported in the Llama-3 paper can be attributed to the key implementation difference: while the Llama-3 paper leverages a comprehensive dataset including a huge amount of Llama-2 model checkpoints to fit their NLL-to-performance curve, our analysis relies on a more limited set of data points. In addition, the compute used to train the sampling LMs also matters for the prediction accuracy. Larger training FLOPs is easier to get emerged performance, which is more useful for fitting the analytical form. The difference in training data density and the FLOPs used to train sampling LMs significantly impacts the sensitivity of the fitted analytical forms, particularly in regions where data points are sparse.
Appendix E FLP-M: Fit Curve for Ablation Study
The FLOPs-to-Loss fit curves are in Fig. 14 and the Loss-to-Performance fit curves are in Fig. 15. M4 in Tab. 3 offers more stable and accurate predictions for domain-specific loss, with the combined approach of domain loss and neural network estimation delivering the best overall downstream performance prediction.
Table 4: The relative prediction error of automated statistical model (Li et al., 2024c), linear function, and neural network in FLP-M Loss-to-Performance mapping.
Method BBH ARC TriviaQA Hellaswag RACE HumanEval
Li et al. (2024c) 14.5 13.2 16.8 4.8 5.9 18.6
Linear Function 4.3 11.8 7.8 2.5 4.1 12.0
Neural Network 9.1 8.6 0.4 1.3 3.4 14.3
Appendix F Automated Statistical Model for FLP-M Loss-to-Performance Mapping
We also conduct experiments with automated statistical model discovery (Li et al., 2024c) to evaluate interpetable approaches (vs. neural networks) for the Loss-to-Performance mapping in FLP-M. Our implementation involves using GPT-4o with multimodal capabilities and 3 maximum iterations. We also apply human post-processing to extract key insights from model-generated natural language feedback. We evaluate the relative performance prediction error across six benchmark tasks (%). Our analysis reveals several key findings:
  • Performance Comparison: The automatically discovered interpretable method shows higher error rates across all benchmarks compared to our neural network approach.
  • Model Complexity Trade-off: While the linear function offers the highest interpretability, it achieves intermediate performance between our neural network and the automated discovery method. This suggests a clear trade-off between model simplicity and prediction accuracy.
  • Neural Network Architecture: Our approach uses a minimalist architecture (two-layer network with hidden size 3 and ReLU activation) that balances interpretability with performance. This simple structure allows us to (a) Express the model in analytical form, (b) Perform detailed numerical analysis, and (c) Maintain transparency in the prediction process
These results demonstrate that while fully interpretable methods are valuable, they currently face challenges in matching the performance of even simple neural networks for this specific task. However, we maintain interpretability in our approach through architectural simplicity and comprehensive analysis capabilities.